OpenStudy (anonymous):

puzzle question : 1_2_3_4_5_6_7_8_9 = 100 fill " _ " with +, -, or * so that LHS=RHS note that bodmas rule doesn't apply here.

OpenStudy (unklerhaukus):

\[(-1+(2\times3)) \times (4\times5)-6+7+8-9 = 100\]

OpenStudy (anonymous):

you used brackets, also theres no _ before 1 .. so i wont take that as the ans @UnkleRhaukus

OpenStudy (unklerhaukus):

bother

OpenStudy (anonymous):

??

OpenStudy (turingtest):

so what are the rules on the order of operations? do you have to do them from left to right, or...?

OpenStudy (anonymous):

i'll show an example 1+2*3*4-5+6+7-8*9=324 you see now??

OpenStudy (turingtest):

so that would be a "yes" you did not do the multiplication first

OpenStudy (turingtest):

and you cannot chose how to arrange the order, you do it from left to right

OpenStudy (unklerhaukus):

i dont get it 1+2*3*4-5+6+7-8*9=39

OpenStudy (turingtest):

he did it like [(1+2)*3*4-5+6+7-8]*9

OpenStudy (unklerhaukus):

*=-39

OpenStudy (turingtest):

left to right...

OpenStudy (turingtest):

[(1+2)*3*4-5+6+7-8]*9 [3*3*4-5+6+7-8]*9 (36)*9=324

OpenStudy (unklerhaukus):

oh

OpenStudy (anonymous):

36*9 = 324 whats the problem? its left to right, where all +,- and * have same priority..

OpenStudy (anonymous):

well not same priority,,but you have to go left to right..

OpenStudy (unklerhaukus):

ah like this? \[((((((((1\star2)\star3)\star4)\star5)\star6)\star7)\star8)\star9) = 100\]

OpenStudy (anonymous):

yep//take it like that..

OpenStudy (turingtest):

division not allowed, right?

OpenStudy (anonymous):

i had my ans without division, i'd like to know a soln without division if you can tell..

OpenStudy (turingtest):

1*2+3*4*5*+6-7-8+9 pretty sure that does the job...

OpenStudy (anonymous):

yep..matches with my soln ! :)

OpenStudy (turingtest):

very cool Q! I've had it for a few minutes, but my browser crashed :(